OpenStudy (anonymous):

Help please! Find the value of k such that -2/3 is a zero of the function f(x) = 4x + k /7. How do you solve this?

OpenStudy (asnaseer):

if -2/3 is a zero this function, then it means that \(f(-2/3)=0\). so just replace \(x\) with \(-2/3\) in the expression for \(f(x)\) and set the while thing equal to zero, then solve to find \(k\).

OpenStudy (asnaseer):

*whole

OpenStudy (anonymous):

\[ f\left(-\frac 23\right)=0\implies 0=4\left(-\frac23\right)+\frac k7 \]This equation allows you to solve for \(k\).

OpenStudy (anonymous):

is it -8/3.... i think i might have done something wrong.

OpenStudy (asnaseer):

please show your steps so that I can spot where you may have made a mistake

OpenStudy (asnaseer):

hang on, is your equation:\[ f(x) = 4x + \frac{k}{7}\tag{1}\]or,\[f(x)=\frac{4x+k}{7}\tag{2}\]

OpenStudy (anonymous):

4 (-2/3) + k /3 = 0. I multiplied 3 on both sides so I got rid of the 3 and made it 4(-2/3) + k = 0. then i solved for 4(-2/3) and ended up with -8/3.

OpenStudy (anonymous):

the second one!

OpenStudy (asnaseer):

you typed: 4 (-2/3) + k /3 = 0. shouldn't that be: (4 (-2/3) + k)/7 = 0.

OpenStudy (anonymous):

oh you're right. my bad haha! let me try that again.

OpenStudy (asnaseer):

you should use brackets to avoid ambiguity, so your question should have been typed as: f(x)=(4x+k)/7

OpenStudy (anonymous):

sorry about that :(

OpenStudy (asnaseer):

np :)

OpenStudy (asnaseer):

in any case, you will still end up with: 4(-2/3) + k = 0 this then leads to: -8/3 + k = 0 you then drew the wrong conclusion from this

OpenStudy (asnaseer):

you need to add 8/3 to both sides of this equation

OpenStudy (anonymous):

so it's positive 8/3, right?

OpenStudy (asnaseer):

yes - well done! :)

OpenStudy (anonymous):

thanks so much! appreciate the help :)

OpenStudy (asnaseer):

yw :)

OpenStudy (anonymous):

thank you @wio too! :)